# Cross-lingual Human-Preference Alignment for Neural Machine Translation with Direct Quality Optimization

### v1

@@ 5.3 Discussion
The nearly-universal [^4] The only language that did not see universal improvements was Hindi, which regressed in BLEU for both datasets, despite improving in the three neural metrics (COMET, CometKiwi, and BLEURT). improvements for both FLORES+ and NTREX in all four automatic translation quality metrics (Table 3) provide strong evidence that DQO is a suitable task-alignment algorithm for the task of producing human-preferred translations.
@@
However, manual inspection of translations before and after DQO revealed language-specific improvements in unrelated languages. In Latvian, for instance, foreign names are transliterated to match Latvian orthography and declined for grammatical case and gender, e.g. Klavinska (2021) report that George Clooney is translated as Džordžs Klūnijs. While the baseline model applies this rule occasionally and inconsistently, the DQO model almost always produces the correct transliteration. Several examples are included in Table 4.
@@
Table 5: Mean number of Multidimensional Quality Metrics (MQM) errors per segment, as annotated by professional human evaluators, with two different groupings: by severity and by whether the MQM subcategory is language specific or agnostic. NT stands for non-translation, i.e., a segment that cannot be construed as a translation of the source. Trivial refers to minor punctuation errors. This covers 100 randomly sampled English segments from the FLORES+ dataset, translated by the NVIDIA Megatron model before task alignment (baseline) and after it (DQO). The weighted MQM score follows Freitag et al. (2021).
@@
One possible explanation is that the baseline model learned to model both transliteration and non-transliteration, due to the range of translation quality in its supervised training data, causing inconsistent behavior at inference time. When DQO then shifts the output distribution towards general high quality behaviors, the probability of any correlated behaviors (e.g., transliteration in Latvian), would also increase.
@@ 5.4 Human Evaluation
To verify the presence of further language-specific changes for unrelated languages, we performed a human evaluation using the Multidimensional Quality Metrics framework (MQM) with professional translators (Lommel et al., 2014; Freitag et al., 2021). The translators were trained on MQM and Anthea [^5] https://github.com/google-research/google-research/tree/a676d87/anthea, the open-source tool we used for performing MQM.
We follow Freitag et al. in weighting major non-translations at 25 MQM points, other major errors at 5, and all minor errors at 1, except minor punctuation errors, which are 0.1 points.
@@
For each language, we sampled complete documents (each generally two to five sentences forming a single paragraph) from FLORES+ until we had 100 source segments, and translated them with the baseline and task-aligned models. These translations were shown with document context (i.e., keeping related segments together) to the translators, who then annotated them.
@@ 6 Related Work
The idea of task–data mismatch in NMT is not new. There has been extensive previous work focused on reducing this mismatch through data filtering, using surface-level heuristics Koehn et al. (2007), statistical and neural alignment and quality evaluation models Sánchez-Cartagena et al. (2018); Heffernan et al. (2022); Peter et al. (2023), language identification models Lui and Baldwin (2011); Joulin et al. (2016), or various combinations of these Koehn et al. (2020).
@@
Perhaps most similar to DQO is DPO MBR fine-tuning, where MBR was used to generate preference pairs for use with DPO Yang et al. (2024). Compared to DQO, this method is more computationally expensive, due to the quadratic costs of MBR, and additionally requires a reference-based QE model. In addition, DQO’s batched online nature ensures that preference pairs remain relevant to the policy model.
@@ 7 Conclusion
We demonstrate the existence of a fundamental task–data mismatch in NMT and introduce Direct Quality Optimization (DQO), an algorithm for aligning a pretrained model with human preference.
Using DQO on a multilingual NMT model, we find improvements in automatic quality metrics for all supported target languages, even those neither used for DQO, nor related to the languages used for DQO. A human evaluation confirms that these improvements also lead to increased human preference.
@@ A.4 Results by Target Language
Model Lang. FLORES+ NTREX
BLEURT COMET CometKiwi BLEU BLEURT COMET CometKiwi BLEU
Baseline bg 0.8510 0.9016 0.8614 42.10 0.7893 0.8592 0.8332 32.85
DQO bg 0.8658 0.9111 0.8708 42.84 0.8063 0.8727 0.8454 33.45
Baseline cs 0.7852 0.8882 0.8413 31.79 0.7365 0.8550 0.8120 30.46
DQO cs 0.8088 0.9069 0.8585 33.12 0.7618 0.8745 0.8325 30.99
Baseline da 0.7748 0.8930 0.8409 45.46 0.7158 0.8554 0.8163 37.71
DQO da 0.7986 0.9094 0.8600 47.94 0.7375 0.8740 0.8364 39.27
Baseline de 0.7514 0.8599 0.8296 39.22 0.6942 0.8204 0.8059 31.32
DQO de 0.7698 0.8731 0.8429 39.71 0.7227 0.8434 0.8243 32.11
Baseline el 0.7407 0.8864 0.8363 27.72 0.7003 0.8694 0.8205 33.05
DQO el 0.7549 0.8934 0.8425 27.95 0.7168 0.8803 0.8259 34.40
Baseline es 0.7522 0.8576 0.8581 27.98 0.7363 0.8500 0.8364 40.87
DQO es 0.7661 0.8670 0.8679 28.84 0.7494 0.8581 0.8471 41.58
Baseline et 0.7848 0.8819 0.8450 26.30 0.7296 0.8461 0.8162 24.25
DQO et 0.8159 0.9024 0.8636 28.21 0.7610 0.8695 0.8398 24.95
Baseline fi 0.8039 0.8927 0.8499 24.28 0.7475 0.8587 0.8290 18.76
DQO fi 0.8364 0.9160 0.8700 26.50 0.7758 0.8793 0.8485 19.59
Baseline fr 0.7737 0.8779 0.8625 51.00 0.6784 0.8332 0.8414 37.09
DQO fr 0.7870 0.8851 0.8679 51.34 0.6950 0.8446 0.8490 38.07
Baseline hi 0.7044 0.7814 0.8167 35.08 0.6494 0.7380 0.7907 26.38
DQO hi 0.7211 0.8032 0.8384 33.96 0.6730 0.7657 0.8193 25.69
Baseline hr 0.8270 0.8989 0.8645 31.37 0.7740 0.8662 0.8354 32.08
DQO hr 0.8407 0.9085 0.8738 32.46 0.7887 0.8780 0.8464 32.29
Baseline hu 0.8565 0.8761 0.8510 27.32 0.7805 0.8241 0.8238 17.76
DQO hu 0.8817 0.8932 0.8671 28.31 0.8050 0.8426 0.8432 18.54
Baseline id 0.8007 0.9068 0.8401 46.16 0.7646 0.8820 0.8107 40.46
DQO id 0.8128 0.9150 0.8503 47.41 0.7789 0.8920 0.8256 41.03
Baseline it 0.7840 0.8788 0.8658 30.32 0.7350 0.8489 0.8321 36.92
DQO it 0.7951 0.8859 0.8725 31.31 0.7557 0.8656 0.8491 37.90
Baseline ja 0.6990 0.8943 0.8589 32.47 0.6221 0.8638 0.8337 26.95
DQO ja 0.7158 0.9058 0.8699 35.00 0.6442 0.8787 0.8511 27.25
Baseline ko 0.6592 0.8714 0.8461 29.63 0.5896 0.8360 0.8144 25.85
DQO ko 0.6876 0.8884 0.8648 30.84 0.6162 0.8565 0.8362 27.43
Baseline lt 0.8084 0.8758 0.8387 26.38 0.7609 0.8452 0.8126 21.93
DQO lt 0.8393 0.8969 0.8577 28.32 0.7867 0.8637 0.8283 22.29
Baseline lv 0.7940 0.8679 0.8269 31.10 0.7066 0.8139 0.7864 20.55
DQO lv 0.8292 0.8906 0.8481 32.95 0.7528 0.8486 0.8172 22.05
Baseline nl 0.7477 0.8619 0.8487 27.70 0.7154 0.8426 0.8254 34.54
DQO nl 0.7665 0.8741 0.8615 28.55 0.7352 0.8606 0.8420 35.88
Baseline no 0.7827 0.8916 0.8561 33.59 0.7447 0.8623 0.8270 36.85
DQO no 0.7963 0.9017 0.8687 34.40 0.7651 0.8782 0.8451 38.87
Baseline pl 0.7728 0.8736 0.8296 21.52 0.7136 0.8389 0.8034 26.32
DQO pl 0.7951 0.8884 0.8421 22.52 0.7356 0.8567 0.8186 27.63
Baseline pt 0.7894 0.8958 0.8480 50.45 0.7090 0.8486 0.8254 34.05
DQO pt 0.7996 0.9008 0.8549 50.53 0.7228 0.8587 0.8355 35.10
Baseline ro 0.8155 0.8995 0.8654 40.57 0.7471 0.8497 0.8346 33.83
DQO ro 0.8298 0.9072 0.8738 41.94 0.7634 0.8637 0.8490 35.53
Baseline ru 0.7616 0.8821 0.8434 32.05 0.6897 0.8408 0.8133 32.85
DQO ru 0.7764 0.8929 0.8541 32.62 0.7087 0.8560 0.8283 33.07
Baseline sl 0.8077 0.8725 0.8428 30.99 0.7250 0.8140 0.7923 28.53
DQO sl 0.8343 0.8913 0.8584 32.32 0.7648 0.8445 0.8209 29.78
Baseline sv 0.7997 0.8970 0.8549 45.30 0.7422 0.8595 0.8203 41.12
DQO sv 0.8141 0.9069 0.8664 46.38 0.7639 0.8784 0.8408 42.41
Baseline tr 0.7821 0.8862 0.8492 29.89 0.6866 0.8271 0.8164 17.59
DQO tr 0.8001 0.8991 0.8631 30.60 0.7116 0.8466 0.8349 17.82
Baseline uk 0.7609 0.8798 0.8318 29.99 0.6892 0.8361 0.7999 25.79
DQO uk 0.7804 0.8930 0.8444 30.93 0.7154 0.8574 0.8177 26.88
Baseline vi 0.7292 0.8775 0.8315 42.13 0.6787 0.8452 0.8096 41.42
DQO vi 0.7488 0.8904 0.8447 43.19 0.6955 0.8594 0.8251 42.24
Baseline zh 0.6948 0.8526 0.8192 40.09 0.6281 0.8106 0.7887 34.54
DQO zh 0.7157 0.8710 0.8359 42.45 0.6509 0.8312 0.8093 36.25
Table 8: Automatic quality evaluation metrics for all target languages supported by the NVIDIA Megatron model, before and after Direct Quality Optimization (DQO), computed on both the FLORES+ and NTREX datasets.

### v2

@@ 5.3 Discussion
The nearly-universal improvements for both FLORES+ and NTREX in all four automatic translation quality metrics (Table 1) provide strong evidence that DQO is a suitable task-alignment algorithm for the task of producing human-preferred translations.
@@
However, manual inspection of translations before and after DQO revealed language-specific improvements in unrelated languages. In Latvian, for instance, foreign names are transliterated to match Latvian orthography and declined for grammatical case and gender, e.g. Klavinska (2021) report that George Clooney should be translated as Džordžs Klūnijs. While the baseline model applies correct transliteration occasionally and inconsistently, the DQO model almost always produces the correct transliteration. Several examples are included in Table 4.
@@
Table 5: Mean number of Multidimensional Quality Metrics (MQM) errors per segment, as annotated by professional human evaluators, with two different groupings: by severity and by whether the MQM subcategory is language specific or agnostic. NT stands for non-translation, i.e., a segment that cannot be construed as a translation of the source. Trivial refers to minor punctuation errors. This covers 100 randomly sampled English segments from the FLORES+ dataset, translated by the NVIDIA Megatron model before task alignment (baseline) and after task alignment (DQO). The weighted MQM score follows Freitag et al. (2021).
@@
One possible explanation is that the baseline model learned to model both transliteration and non-transliteration, due to the range of translation quality in its supervised training data, causing inconsistent behavior at inference time. When DQO then shifts the output distribution towards certain human-preferred features, the probability of any correlated features (e.g., transliteration in Latvian), also increases.
@@ 5.4 Human Evaluation
To verify the presence of further language-specific changes for unrelated languages, we performed a human evaluation using the Multidimensional Quality Metrics framework (MQM) with professional translators (Lommel et al., 2014; Freitag et al., 2021). The translators were trained on MQM and Anthea [^4] https://github.com/google-research/google-research/tree/a676d87/anthea, the open-source tool we used for performing MQM. We follow Freitag et al. in weighting major non-translations at 25 MQM points, other major errors at 5, and all minor errors at 1, except minor punctuation errors, which are 0.1 points.
@@
For each language, we sampled complete documents (each generally two to five sentences forming a single paragraph) from FLORES+ until we had 100 source segments. The translators then annotated the baseline and task-aligned translations.
@@ 6 Related Work
The idea of task–data mismatch in NMT is not new. There has been extensive previous work focused on reducing this mismatch through data filtering, using surface-level heuristics Koehn et al. (2007), statistical and neural models for alignment and quality evaluation Sánchez-Cartagena et al. (2018); Heffernan et al. (2022); Peter et al. (2023), language identification Lui and Baldwin (2011); Joulin et al. (2016), or ensembles Koehn et al. (2020).
@@
In DPO MBR fine-tuning, MBR is used to generate preference pairs for use with DPO Yang et al. (2024). Compared to DQO, this method is computationally more expensive, and requires a reference-based QE model. In addition, DQO’s online nature ensures that preference pairs remain relevant to the policy model.
Reward rAnked Fine-Tuning (RAFT) is the method most similar to DQO, but uses SFT to update the model towards a single preferred output rather than using DPO with a preferred/rejected output pair Dong et al. (2023). As it was not evaluated for the translation task, used an independently trained reward model, and had slight differences in sampling parameters, we ran an ablation on whether to use DPO or SFT in DQO (see Section 4).
@@ 7 Conclusion
We demonstrate the existence of a fundamental task–data mismatch in NMT and introduce Direct Quality Optimization (DQO), a method of aligning pretrained models with human preference.
Using DQO on a multilingual NMT model, we find improvements in automatic quality metrics for all supported target languages, even those neither used for DQO, nor related to the languages used for DQO. A human evaluation confirms that these improvements reflect increased human preference.
@@
8 Limitations
This work only tests one quality evaluation model as a proxy for human preferences, CometKiwi, and does not examine the impact of that proxy’s quality. We focused primarily on a single translation model, the NVIDIA Megatron English-Many model, using a 1.3B paramter English-German model only for the perplexity experiments (as we had access to the training data). Human evaluation of translation quality was only performed on two language pairs. For all others, we relied on automatic quality evaluation metrics such as BLEURT, COMET and BLEU, which may not fully capture true human preference.
@@ A.4 Results by Target Language
Model Lang. FLORES+ devtest NTREX
BLEURT COMET22 CometKiwi22 BLEU BLEURT COMET22 CometKiwi22 BLEU
Baseline bg 0.8400 0.8974 0.8524 41.80 0.7713 0.8520 0.8242 32.00
DQO bg 0.8526 0.9067 0.8614 42.70 0.7865 0.8638 0.8341 32.40
Baseline cs 0.7758 0.8826 0.8327 32.60 0.7282 0.8509 0.8065 30.10
DQO cs 0.7978 0.9002 0.8504 34.00 0.7506 0.8696 0.8255 30.70
Baseline da 0.7744 0.8942 0.8396 46.40 0.7136 0.8541 0.8145 37.40
DQO da 0.7948 0.9091 0.8565 48.60 0.7355 0.8721 0.8341 39.30
Baseline de 0.7417 0.8535 0.8222 38.80 0.6793 0.8100 0.7950 30.80
DQO de 0.7561 0.8682 0.8338 39.30 0.7041 0.8315 0.8117 31.80
Baseline el 0.6738 0.8641 0.8032 25.90 0.6477 0.8494 0.7876 30.60
DQO el 0.6793 0.8699 0.8044 26.60 0.6567 0.8585 0.7892 31.60
Baseline es 0.7467 0.8567 0.8569 27.50 0.7304 0.8474 0.8330 40.50
DQO es 0.7594 0.8656 0.8662 28.80 0.7421 0.8547 0.8425 41.00
Baseline et 0.7779 0.8792 0.8421 27.10 0.7279 0.8451 0.8155 24.20
DQO et 0.8114 0.9041 0.8647 28.90 0.7603 0.8690 0.8399 25.00
Baseline fi 0.7959 0.8899 0.8471 24.40 0.7393 0.8550 0.8247 18.70
DQO fi 0.8264 0.9105 0.8640 26.00 0.7640 0.8736 0.8421 19.60
Baseline fr 0.7400 0.8638 0.8486 49.40 0.6525 0.8221 0.8289 36.10
DQO fr 0.7529 0.8713 0.8544 50.70 0.6632 0.8305 0.8344 37.00
Baseline hi 0.6825 0.7645 0.8040 32.90 0.6313 0.7227 0.7735 25.50
DQO hi 0.6991 0.7862 0.8217 32.50 0.6511 0.7459 0.7972 25.10
Baseline hr 0.8190 0.8942 0.8624 31.10 0.7707 0.8644 0.8326 31.80
DQO hr 0.8318 0.9032 0.8695 32.10 0.7847 0.8770 0.8445 32.50
Baseline hu 0.8378 0.8645 0.8354 26.90 0.7616 0.8141 0.8118 17.40
DQO hu 0.8554 0.8800 0.8488 27.10 0.7793 0.8294 0.8268 18.00
Baseline id 0.8030 0.9092 0.8414 47.50 0.7648 0.8823 0.8111 40.50
DQO id 0.8158 0.9172 0.8516 49.30 0.7784 0.8917 0.8251 41.10
Baseline it 0.7699 0.8725 0.8590 30.60 0.7280 0.8455 0.8279 36.70
DQO it 0.7860 0.8821 0.8676 31.40 0.7467 0.8613 0.8434 37.50
Baseline ja 0.6832 0.8918 0.8545 32.60 0.6042 0.8584 0.8251 26.40
DQO ja 0.6981 0.9019 0.8629 34.10 0.6208 0.8713 0.8395 27.10
Baseline ko 0.6538 0.8689 0.8433 29.40 0.5788 0.8317 0.8085 25.50
DQO ko 0.6734 0.8820 0.8550 30.30 0.5980 0.8481 0.8250 26.50
Baseline lt 0.8043 0.8742 0.8344 27.30 0.7485 0.8404 0.8057 21.60
DQO lt 0.8264 0.8910 0.8490 28.80 0.7699 0.8564 0.8181 22.30
Baseline lv 0.7896 0.8677 0.8253 30.50 0.6997 0.8097 0.7816 20.40
DQO lv 0.8201 0.8902 0.8431 32.10 0.7418 0.8424 0.8088 21.70
Baseline nl 0.7425 0.8617 0.8483 27.00 0.7080 0.8384 0.8205 34.20
DQO nl 0.7611 0.8756 0.8601 28.10 0.7262 0.8556 0.8356 35.40
Baseline no 0.7771 0.8899 0.8526 33.80 0.7447 0.8622 0.8267 36.90
DQO no 0.7915 0.8991 0.8646 34.00 0.7644 0.8779 0.8445 38.70
Baseline pl 0.7600 0.8678 0.8206 21.40 0.6992 0.8312 0.7939 25.70
DQO pl 0.7787 0.8818 0.8312 22.80 0.7153 0.8463 0.8058 26.80
Baseline pt 0.7856 0.8941 0.8453 50.80 0.7069 0.8477 0.8236 33.90
DQO pt 0.7952 0.9000 0.8531 51.20 0.7197 0.8574 0.8341 35.00
Baseline ro 0.8026 0.8927 0.8594 40.30 0.7338 0.8441 0.8255 33.30
DQO ro 0.8144 0.9015 0.8645 41.40 0.7474 0.8571 0.8386 34.70
Baseline ru 0.7430 0.8755 0.8329 31.30 0.6706 0.8299 0.8002 31.80
DQO ru 0.7556 0.8842 0.8419 32.00 0.6831 0.8433 0.8104 31.90
Baseline sl 0.7978 0.8679 0.8359 30.00 0.7174 0.8106 0.7877 28.30
DQO sl 0.8252 0.8860 0.8517 31.80 0.7576 0.8410 0.8163 29.60
Baseline sv 0.7945 0.8957 0.8515 45.40 0.7401 0.8581 0.8192 40.90
DQO sv 0.8113 0.9064 0.8650 46.20 0.7632 0.8781 0.8400 42.40
Baseline tr 0.7693 0.8827 0.8441 29.10 0.6802 0.8235 0.8129 17.60
DQO tr 0.7875 0.8953 0.8559 30.10 0.7011 0.8402 0.8287 17.70
Baseline uk 0.7432 0.8728 0.8172 29.80 0.6678 0.8230 0.7838 24.80
DQO uk 0.7603 0.8878 0.8300 30.50 0.6868 0.8423 0.7983 25.80
Baseline vi 0.7157 0.8736 0.8299 42.20 0.6753 0.8442 0.8081 41.30
DQO vi 0.7329 0.8857 0.8429 43.80 0.6917 0.8589 0.8234 42.10
Baseline zh 0.7015 0.8582 0.8199 42.00 0.6267 0.8099 0.7879 34.50
DQO zh 0.7202 0.8752 0.8367 44.10 0.6468 0.8292 0.8067 36.00
Table 8: Automatic quality evaluation metrics for all target languages supported by the NVIDIA Megatron model, before and after Direct Quality Optimization (DQO), computed on both the FLORES+ devtest and NTREX datasets.
A.5 Ablation of Update Step: DPO vs. SFT
Figure 4: Mean BLEURT-20 on FLORES+ dev after each round of DQO with the NVIDIA Megatron EN-X model, using either Direct Preference Optimization (DPO) or Supervised Fine-Tuning (SFT) to update the model. DQO with SFT is equivalent to Reward rAnked Fine-Tuning (RAFT).